# PADE: A Predictor-Free Sparse Attention Accelerator via Unified Execution and Stage Fusion

### v1

@@ I Introduction
Transformer models have achieved significant success in various fields, spanning from content generation [song2018situ, brown2020language, hendrycks2020measuring, lin2023videodirectorgpt] to computer vision [dosovitskiy2020image, you2023vitcod, bai2024seed]. However, the self-attention mechanism used in Transformers suffers from quadratic time and memory complexity, limiting their scalability to long sequences.
To address this, sparse attention techniques have emerged as a promising solution, where attention is computed over a subset of query-key (Q-K) pairs instead of a dense attention matrix. Existing sparse attention works can be divided into two routes: static sparsity (SS) and dynamic sparsity (DS). SS [parmar2018image, child2019generating, qiu2019blockwise, zaheer2020big, beltagy2020longformer, jiang2024minference, dong2022cswin, katharopoulos2020transformers] relies on predefined sparse patterns that remain fixed during inference, lacking flexibility and often resulting in significant accuracy degradation [li2024large, fuad2023survey, kachris2025survey]. In contrast, DS [correia2019adaptively, kitaev2020reformer, you2023vitcod, wang2023cta, song2024tsacc, cho2024sparc, park2024token, fan2022adaptable, dass2023vitality, li2020ftrans, wang2024sofa, qin2023fact, ham20203, ham2021elsa, lu2021sanger, qu2022dota, yang2022dtatrans, zhou2022energon, wang2025bitstopper, wang2021spatten, liu2022dynamic, wang2025lapa, li2022accelerating, wang2025mcbp] adapts the sparsity pattern at runtime, offering improved accuracy and flexibility, making it more suitable for a broader range of tasks and input types.
@@
Unfortunately, such an added predictor occupies substantial overhead, increasingly offsetting the sparsity benefit. Fig. 2 (a) shows the power breakdown of dense attention and two representative DS accelerators: Sanger [lu2021sanger], SOFA [wang2024sofa], with varying executor bit-widths. Power consumption is categorized into executor and predictor components. Sanger uses 4-bit MSB multiplication with threshold comparison, while SOFA employs log domain shifting with top-$k$ sorting. As depicted in Fig. 2(a), there are two key observations: (1) At larger executor bit-widths (e.g., 16bit), the DS reduces overall power by about $63\%$, with the predictor costing only $33\%$. This explains why previous DS works adopt additional sparsity predictors. (2) However, as executor bit-width decreases, predictor overhead becomes dominant. At 8-bit, overall savings drop to merely $32\%$, with the sparsity predictor occupying over $63\%$ of total cost. This is due to the predictor must access and process full-sized K tensors, a cost unaffected by sparsity.
@@
Takeaway: With the rapid advancement of Transformer quantization techniques, like GPTQ [frantar2022gptq], LLM.int8() [dettmers2022gpt3], SmoothQuant [xiao2023smoothquant] and Atom [zhao2024atom], there is a growing trend of adopting low-bit quantization in attention mechanisms. In these cases, the predictor’s overhead increasingly offsets the benefits of sparsity. This highlights a need to reduce or even eliminate the prediction overhead.
Insights: The root cause of the excessive prediction cost stems from the decoupling between existing sparsity predictors and executors, which hinders the computational and memory access efforts paid in the predictor from being reused by the executor. To address this, we draw inspiration from bit-serial computing [judd2016stripes, albericio2017bit, lee2018unpu, chen2024bbs, kam2025panacea, im2023sibia, guo2025transitive], which separates an INT operation into multi-round bit-level steps. This motivates a unified design that integrates prediction and execution into a single computation stage, thereby eliminating the separate predictor and improving overall efficiency.
@@ III-A Re-examinating DS Works and Opportunity
As shown in Fig. 4(a), traditional DS works adopt a stage-splitting paradigm that decouples prediction from execution. Taking Sanger [lu2021sanger] as an example, the predictor takes the full 4-bit Key tensor to identify iQKs, while the executor separately fetches the corresponding KVs for precise computation. However, this paradigm introduces two inefficiency sources: (1) For Keys related to iEQKs (e.g., K0), a single bit may suffice to identify their insignificance. However, the stage-splitting approach blindly loads 4 bits, resulting in redundancy. (2) For Keys related to iQKs (e.g., Ks), the executor reloads their high-bit-width versions for subsequent more precise computation. However, it fails to reuse the data already processed during prediction, leading to inefficiency.
@@
Unlike stage-splitting DS designs that rely on separate predictors, BSF eliminates prediction overhead and enables fine-grained early termination to reduce unnecessary computation and memory access. Fig. 4(c) profiles the memory access and computation reduction achieved by different strategies in the attention modules across four layers from LLaMA-2 7B [touvron2023llama2]. On average, BSF can achieve $4.6\times$ higher memory access and $2.1\times$ more computation reduction, compared to traditional stage-splitting DS approaches, highlighting its significant energy efficiency advantage over current DS methods.
@@ III-B Challenges for Bit-serial Enable Stage-fusion
Key idea We introduce the concept of a bit uncertainty interval (BUI), which defines the maximum fluctuation range of dot products across diverse bit-planes, via simple yet efficient bit-flipping. To support this, we design a scoreboard-based result-reuse PE to reduce BUI hardware overhead and redundant bit-plane memory accesses.
@@
As shown in Fig. 5 (d), to enable bit-grained early termination, it is critical to avoid the bulk loading of all bit-planes for each Key. Instead, each Key should be evaluated bit-plane by bit-plane to determine whether it qualifies as an iEQK. If true, the subsequent bit-planes for that Key are skipped; otherwise, the next bit-plane is required. However, due to DRAM’s dynamic precharge mechanism, loading each bit-plane typically incurs several dozen cycles [chang2016understanding, JEDEC_JESD235D]. Naively stalling computation during data loading results in underutilized computational resources.
Key idea. We propose bidirectional sparsity (BS), treating bit ‘1’ as a form of sparsity, thus ensuring load imbalance is below $50\%$. Building upon this, we introduce bit-wise out-of-order (OOE) execution, allowing the PE to process other bit-planes while avoiding memory access stalls. To support this, we propose a lightweight ANDer tree to reduce multiplexer’s overhead from BS sparsity, while the scoreboard-based PE facilitates partial sum buffering and reuse.
@@
As shown in Fig. 5 (e), existing pruning strategies rely on row-wise attention score distributions to assess token importance, introducing strong row dependencies. This dependency prevents effective tiling, which is essential for IO efficiency. Without tiling, when the number of parallel queries increases, memory access overhead grows sharply. As illustrated in Fig. 5 (f), increasing the number of parallel queries ($P$) from 8 to 32, leads to over 12× more memory accesses. A coarse solution is to enlarge on-chip SRAM, but this incurs significant area inefficiency. For example, with (PP=512, SS=2048), 5MB of SRAM is required, resulting in a 5.47 mm2 footprint under TSMC 28nm technology, which is $7.4\times$ and $8.9\times$ larger than the total area of SpAtten [wang2021spatten] and ELSA [ham2021elsa], respectively.
@@
ELSA [ham2021elsa] $\times$ $\times$ $\times$ Value
Sanger [lu2021sanger] $\times$ $\times$ $\times$ Value
DOTA [qu2022dota] $\times$ $\times$ $\times$ Value
DTATrans[yang2022dtatrans] Low ✓* $\times$ Value
SpAtten [wang2021spatten] Low ✓* $\times$ Multi-bit
Energon [zhou2022energon] $\times$ $\times$ $\times$ Multi-bit
FACT [qin2023fact] $\times$ $\times$ $\times$ Value
SOFA [wang2024sofa] Low $\times$ Value
@@
Unfortunately, current attention accelerators still suffer from computation and memory access inefficiencies, as they fail to exploit bit-grained opportunities to eliminate the high-overhead predictor. Table I summarizes their features. The majority of existing works [ham20203, ham2021elsa, lu2021sanger, qu2022dota, yang2022dtatrans] focus on accelerating attention by alleviating computation overhead. For example, ELSA [ham2021elsa], Sanger [lu2021sanger], DOTA [qu2022dota], FACT [qin2023fact] adopt techniques like binary hashing, half-precision MSB, low-rank approximation, log-domain shifting to accelerate computation. However, these methods overlook memory optimization. While SpAtten [wang2021spatten]. SOFA [wang2024sofa] realizes this challenge, their coarse-grained strategies, like hybrid quantization, and cross-stage tiling, fail to exploit fine-grained, bit-level optimizations. Further, all current works rely on extra sparsity predictors, incurring substantial overhead. Notably, DTATrans [yang2022dtatrans] and SpAtten [wang2021spatten] guide sparsity using attention scores from the previous layer. While this strategy partially reduces predictor overhead, it necessitates resource-intensive retrain to recover accuracy. These limitations motivate us to design an efficient attention accelerator that jointly optimizes computation and memory at fine granularity, while eliminating the sparsity predictor.
@@ IV-B BS-OOE: Improving Hardware Utilization
The core idea of BS is that 1 is also a form of sparsity. Without loss of generality, the dot-product between an $N$-element Query and Key can be formulated as:
@@ IV-C ISTA: Enhancing IO Efficiency
However, a key observation is that naive left-to-right computation causes frequent updates of the maximum across tiles (line 9 in Fig. 10 (c)), leading to a series of redundant operations (line 11). Specifically, each time the maximum is updated, it triggers one subtraction, one exponentiation and two multiplications of scalar and vector (see highlights in lines 11-12). We employ the arithmetic complexity model [brent2010modern, wang2021efficientMIMO, wang2023efficient] to normalize the complexity for different operations. As profiled in Fig. 10 (b), for SS=$2048$ and $B_{c}$=16, this leads to about 30% increase in computation complexity compared to the vanilla implementation. Moreover, the overhead becomes more as $B_{c}$ decreases, due to more frequent updates.
To minimize the overhead associated with the ‘max updating’, it is advantageous to prioritize dominant tokens. Note that the reduction in operations refers to the redundant computations shown in lines 11–12, rather than the comparison operations in line 9. To this end, we propose a head-tail interleaved updating strategy that exploits the locality property of attention: Recently generated tokens and the initial token typically exhibit higher weights than others [tian2023scan, jiang2024minference]. As illustrated in Fig. 10(a), the update begins with the initial region, then jumps to the recent region, and subsequently returns to the post-initial region, repeating this interleaved pattern across tiles. This pattern reduces unnecessary maximum value updates, resulting in a $20\%$-$40\%$ reduction in total operations. It is important to note that, without attention locality, the performance of head-tail interleaving is on par with regular execution and not worse.
@@ V-D Grouped Lightweight Sparsity ANDer Tree (GSAT)
BS Scheduler. PADE adopts a low-cost BS scheduler to manage the operations within each PE, as depicted in Fig. 12. To control the bit-serial dot product, the scheduler first identifies whether a bit plane of the Key vector contains more zeros or ones. It then sends the original or flipped bit column to a priority encoder. The priority encoder operates on five consecutive bits of the plane at a time. For example, at the first time, it receives $k_{0}$$k_{4}$}, followed by $k_{1}$$k_{5}$} in the next, and so on. The encoder detects the location of the first “1” bit within each 5-bit vector. If such a bit exists, it is masked, and the remaining bits are propagated to the next time step. Otherwise, if the vector contains only zeros, the encoder asserts val=0 to disable the corresponding bit-serial multiplier in the PE, as depicted in Fig. 11 (b).
Figure 12: Detailed architecture of bidirectional sparsity (BS) scheduler.
V-E Reuse-Aware Reorder Scheduling (RARS)
VI Evaluation
VI-A Experimental Setup
Baseline comparisons: We compare PADE with five SOTA attention accelerators: Sanger [lu2021sanger], Spatten [wang2021spatten], Energon [zhou2022energon], DOTA [qu2022dota], SOFA [wang2024sofa]. For fair comparison, all designs are normalized to a 28nm process and evaluated under identical conditions: PE arrays occupy the same area as PADE and work in 800 MHz, on-chip SRAM is set to 352kB, and peak HBM bandwidth is 256 GB/s, with 4 pj/bit [o2017fine].
@@ VI-A Experimental Setup
Benchmarks: We evaluate PADE on several representative Transformer models across NLP and CV tasks. For NLP tasks, Llama2-7B [touvron2023llama2], Llama3-8B [grattafiori2024llama], Qwen7B [bai2023qwen], Bloom1B7 [le2022bloom] and OPT1B3 [zhang2022opt], for six tasks.These tasks include language modeling (Wikitext-2 (S=2k) [merity2016pointer], Wikilingua (S=2k) [ladhak-wiki-2020], Winogrande (S=0.25k) [sakaguchi2021winogrande]), language understanding (MMLU, S=0.5k) [hendrycks2020measuring], code generation MBPP (S=1k) [austin2021program], long context processing dolly (S=15k) [conover2023free]. For CV tasks, we choose ViT-L/16 (S=576) [dosovitskiy2020image] and PVT (S=3k) [wang2021pyramid] on ImageNet-1k [deng2009imagenet] and JFT [sun2017revisiting] classification.
Quantization Accuracy. All pre-trained models are sourced from Pytorch [paszke2017automatic] and HuggingFace [wolf2020transformers]. INT8 baselines derived via post-training quantization, where only the weights and activations (QKVs) are quantized to INT8, while non-linear operators (e.g., softmax) remain in FP16 precision. As shown in Table. II, the INT8 baseline incurs less than a $1\%$ average accuracy drop from FP16, thus confirming its validity.
Hardware Evaluation: Table III lists the hardware configuration of PADE. We implement the RTL design for PADE and utilize Synopsys DC on TSMC 28nm CMOS technology to estimate the logic area and power. The power, area, and read/write bandwidth of on-chip SRAM buffers are estimated through CACTI [muralimanohar2009cacti]. Off-chip HBM modeling involves simulating access patterns and row activation under various data layouts (Fig. 22), capturing HBM’s burst behavior. We derive memory latency from Ramulator [kim2015ramulator], and estimate IO power following the methodology in [cavigelli2016origami, andri2016yodann, wang2017energy]. We extract each stage’s cycles by simulating the RTL with Verilator [snyder2004verilator], and use a custom cycle-level simulator to evaluate performance.

### v2

@@ I Introduction
Transformer models have achieved significant success in various fields, spanning from content generation [107, 5, 47, 75] to computer vision [27, 145, 7]. However, the self-attention mechanism used in Transformers suffers from quadratic time and memory complexity, limiting their scalability to long sequences.
To address this, sparse attention techniques have emerged as a promising solution, where attention is computed over a subset of query-key (Q-K) pairs instead of a dense attention matrix. Existing sparse attention works can be divided into two routes: static sparsity (SS) and dynamic sparsity (DS). SS [89, 16, 94, 146, 9, 56, 26, 61] relies on predefined sparse patterns that remain fixed during inference, lacking flexibility and often resulting in significant accuracy degradation [72, 34, 58]. In contrast, DS [19, 63, 145, 117, 108, 17, 88, 29, 20, 69, 119, 93, 41, 42, 81, 95, 142, 150, 121, 116, 76, 122, 74, 125] adapts the sparsity pattern at runtime, offering improved accuracy and flexibility, making it more suitable for a broader range of tasks and input types.
@@
Unfortunately, such an added predictor occupies substantial overhead, increasingly offsetting the sparsity benefit. Fig. 2 (a) shows the power breakdown of dense attention and two representative DS accelerators: Sanger [81], SOFA [119], with varying executor bit-widths. Power consumption is categorized into executor and predictor components. Sanger uses 4-bit MSB multiplication with threshold comparison, while SOFA employs log domain shifting with top-$k$ sorting. As depicted in Fig. 2(a), there are two key observations: (1) At larger executor bit-widths (e.g., 16bit), the DS reduces overall power by about $63\%$, with the predictor costing only $33\%$. This explains why previous DS works adopt additional sparsity predictors. (2) However, as executor bit-width decreases, predictor overhead becomes dominant. At 8-bit, overall savings drop to merely $32\%$, with the sparsity predictor occupying over $63\%$ of total cost. This is due to the predictor must access and process full-sized K tensors, a cost unaffected by sparsity.
@@
Takeaway: With the rapid advancement of Transformer quantization techniques, like GPTQ [33], LLM.int8() [25], SmoothQuant [137] and Atom [149], there is a growing trend of adopting low-bit quantization in attention mechanisms. In these cases, the predictor’s overhead increasingly offsets the benefits of sparsity. This highlights a need to reduce or even eliminate the prediction overhead.
Insights: The root cause of the excessive prediction cost stems from the decoupling between existing sparsity predictors and executors, which hinders the computational and memory access efforts paid in the predictor from being reused by the executor. To address this, we draw inspiration from bit-serial computing [57, 1, 66, 15, 59, 52, 40], which separates an INT operation into multi-round bit-level steps. This motivates a unified design that integrates prediction and execution into a single computation stage, thereby eliminating the separate predictor and improving overall efficiency.
@@ III-A Re-examinating DS Works and Opportunity
As shown in Fig. 4(a), traditional DS works adopt a stage-splitting paradigm that decouples prediction from execution. Taking Sanger [81] as an example, the predictor takes the full 4-bit Key tensor to identify iQKs, while the executor separately fetches the corresponding KVs for precise computation. However, this paradigm introduces two inefficiency sources: (1) For Keys related to iEQKs (e.g., K0), a single bit may suffice to identify their insignificance. However, the stage-splitting approach blindly loads 4 bits, resulting in redundancy. (2) For Keys related to iQKs (e.g., Ks), the executor reloads their high-bit-width versions for subsequent more precise computation. However, it fails to reuse the data already processed during prediction, leading to inefficiency.
@@
Unlike stage-splitting DS designs that rely on separate predictors, BSF eliminates prediction overhead and enables fine-grained early termination to reduce unnecessary computation and memory access. Fig. 4(c) profiles the memory access and computation reduction achieved by different strategies in the attention modules across four layers from LLaMA-2 7B [113]. On average, BSF can achieve $4.6\times$ higher memory access and $2.1\times$ more computation reduction, compared to traditional stage-splitting DS approaches, highlighting its significant energy efficiency advantage over current DS methods.
@@ III-B Challenges for Bit-serial Enable Stage-fusion
Key idea. Inspired the conservative margin concept [74, 88], we propose bit uncertainty interval-enabled guarded filtering (BUI-GF), a lightweight mechanism for simple yet accurate max-based pruning. BUI-GF characterizes the maximum possible fluctuation of dot products across bit planes using simple yet efficient bit flipping, enabling conservative yet hardware-efficient pruning decisions with minimal overhead. To support this, we design a scoreboard-based result-reuse PE to reduce BUI hardware overhead and redundant bit-plane memory accesses.
@@
As shown in Fig. 5 (d), to enable bit-grained early termination, it is critical to avoid the bulk loading of all bit-planes for each Key. Instead, each Key should be evaluated bit-plane by bit-plane to determine whether it qualifies as an iEQK. If true, the subsequent bit-planes for that Key are skipped; otherwise, the next bit-plane is required. However, due to DRAM’s dynamic precharge mechanism, loading each bit-plane typically incurs several dozen cycles [12, 55]. Naively stalling computation during data loading results in underutilized computational resources.
Key idea. Inspired by bidirectional sparsity [15] and XNOR-BNN-based formulations [35], we propose a two-pronged approach to mitigate resource underutilization. First, we introduce bidirectional, runtime-adaptive sparsity orchestration for the K matrix, which dynamically interprets bit ‘1’ as sparsity in coordination with queries, ensuring load imbalance remains below $50\%$. Building upon this, we further introduce bit-wise out-of-order (OOE) execution, allowing the PE to process other bit-planes while avoiding memory access stalls. To support this, we propose a temporal-reuse-based sparsity scheduler to alleviate runtime scheduling overhead, a lightweight ANDer tree to BS-induced multiplexing overhead, and a scoreboard-based PE that facilitates partial-sum buffering and reuse.
@@
As shown in Fig. 5 (e), existing pruning strategies rely on row-wise attention score distributions to assess token importance, introducing strong row dependencies. This dependency prevents effective tiling, which is essential for IO efficiency. Without tiling, when the number of parallel queries increases, memory access overhead grows sharply. As illustrated in Fig. 5 (f), increasing the number of parallel queries ($P$) from 8 to 32, leads to over 12× more memory accesses. A coarse solution is to enlarge on-chip SRAM, but this incurs significant area inefficiency. For example, with (PP=512, SS=2048), 5MB of SRAM is required, resulting in a 5.47 mm2 footprint under TSMC 28nm technology, which is $7.4\times$ and $8.9\times$ larger than the total area of SpAtten [116] and ELSA [42], respectively.
@@
ELSA [42] $\times$ $\times$ $\times$ Value
Sanger [81] $\times$ $\times$ $\times$ Value
DOTA [95] $\times$ $\times$ $\times$ Value
DTATrans[142] Low ✓* $\times$ Value
SpAtten [116] Low ✓* $\times$ Multi-bit
Energon [150] $\times$ $\times$ $\times$ Multi-bit
FACT [93] $\times$ $\times$ $\times$ Value
SOFA [119] Low $\times$ Value
@@
Unfortunately, current attention accelerators still suffer from computation and memory access inefficiencies, as they fail to exploit bit-grained opportunities to eliminate the high-overhead predictor. Table I summarizes their features. The majority of existing works [41, 42, 81, 95, 142] focus on accelerating attention by alleviating computation overhead. For example, ELSA [42], Sanger [81], DOTA [95], FACT [93] adopt techniques like binary hashing, half-precision MSB, low-rank approximation, log-domain shifting to accelerate computation. However, these methods overlook memory optimization. While SpAtten [116]. SOFA [119] realizes this challenge, their coarse-grained strategies, like hybrid quantization, and cross-stage tiling, fail to exploit fine-grained, bit-level optimizations. Further, all current works rely on extra sparsity predictors, incurring substantial overhead. Notably, DTATrans [142] and SpAtten [116] guide sparsity using attention scores from the previous layer. While this strategy partially reduces predictor overhead, it necessitates resource-intensive retrain to recover accuracy. These limitations motivate us to design an efficient attention accelerator that jointly optimizes computation and memory at fine granularity, while eliminating the sparsity predictor.
@@ IV-B BS-OOE: Improving Hardware Utilization
We extend the BS from static weight scenarios [15] to dynamic attention workloads. The core idea of BS is that bit value ‘1’ is also a form of sparsity. Without loss of generality, the dot-product between an $N$-element Query and Key can be formulated as:
@@
Notably, unlike prior work [15] that targets static weight compression or bit skipping, we extend bidirectional sparsity as a runtime load-balancing mechanism to bound PE imbalance in bit-serial QK execution, where a lightweight sparsity-aware scheduler (see §V-D) is crucial due to the highly dynamic, runtime-determined nature of attention workloads.
@@ IV-C ISTA: Enhancing IO Efficiency
However, a key observation is that naive left-to-right computation causes frequent updates of the maximum across tiles (line 9 in Fig. 10 (c)), leading to a series of redundant operations (line 11). Specifically, each time the maximum is updated, it triggers one subtraction, one exponentiation and two multiplications of scalar and vector (see highlights in lines 11-12). We employ the arithmetic complexity model [10, 120, 118] to normalize the complexity for different operations. As profiled in Fig. 10 (b), for SS=$2048$ and $B_{c}$=16, this leads to about 30% increase in computation complexity compared to the vanilla implementation. Moreover, the overhead becomes more as $B_{c}$ decreases, due to more frequent updates.
To minimize the overhead associated with the ‘max updating’, it is advantageous to prioritize dominant tokens. Note that the reduction in operations refers to the redundant computations shown in lines 11–12, rather than the comparison operations in line 9. To this end, we propose a head-tail interleaved updating strategy that exploits the locality property of attention: Recently generated tokens and the initial token typically exhibit higher weights than others [112, 56]. As illustrated in Fig. 10(a), the update begins with the initial region, then jumps to the recent region, and subsequently returns to the post-initial region, repeating this interleaved pattern across tiles. This pattern reduces unnecessary maximum value updates, resulting in a $20\%$-$40\%$ reduction in total operations. It is important to note that, without attention locality, the performance of head-tail interleaving is on par with regular execution and not worse.
@@ V-D Grouped Lightweight Sparsity ANDer Tree (GSAT)
Figure 12: Temporal reuse architecture of bidirectional sparsity (BS) scheduler.
@@
Lightweight BS Scheduler. Built upon [15], PADE adopts a low-cost BS scheduler by reusing a critical priority encoder across time steps to orchestrate operations within each PE, as depicted in Fig. 12. Unlike the straightforward design in [15], which instantiates multiple priority encoders in parallel, PADE temporally multiplexes a single priority encoder across successive time steps for index selection. This temporal reuse is enabled by PADE’s distinctive pipelined microarchitecture, where the QK-PU and V-PU operate in a staggered pipeline, effectively hiding the additional latency introduced by encoder reuse. The temporal reuse scheme helps PADE reduce 75% priority encoder overhead.
To control the bit-serial dot product, the scheduler first identifies whether a bit plane of the Key vector contains more zeros or ones. It then sends the original or flipped bit column to a priority encoder. The priority encoder operates on five consecutive bits of the plane at a time. For example, at the first time, it receives $k_{0}$$k_{4}$}, followed by $k_{1}$$k_{5}$} in the next, and so on. The encoder detects the location of the first “1” bit within each 5-bit vector. If such a bit exists, it is masked, and the remaining bits are propagated to the next time step. Otherwise, if the vector contains only zeros, the encoder asserts Vx=0 to disable the corresponding bit-serial multiplier in the PE, as depicted in Fig. 11 (b).
@@
V-E Reuse-Aware Reorder Scheduling (RARS)
VI Evaluation
VI-A Experimental Setup
Baseline comparisons: We compare PADE with five SOTA attention accelerators: Sanger [81], Spatten [116], Energon [150], DOTA [95], SOFA [119]. For fair comparison, all designs are normalized to a 28nm process and evaluated under identical conditions: PE arrays occupy the same area as PADE and work in 800 MHz, on-chip SRAM is set to 352kB, and peak HBM bandwidth is 256 GB/s, with 4 pj/bit [86].